OpenStudy (anonymous):

(2^4)^3 can someone help me

OpenStudy (gw2011):

Do you know how to calculate an exponent raised to an exponent? Taking an easier example, do you know how to calculate a number raised to anexponent?

OpenStudy (anonymous):

can you give me a example

OpenStudy (anonymous):

its 2 to the 4th power then that number to the 3rd power right?

OpenStudy (gw2011):

For example, how would you calculate 3^2?

OpenStudy (anonymous):

yes kitty

OpenStudy (anonymous):

ok then you would do 2 times 2 which is 4 then times 2 again which is 8 then one more time making it a forth time times 2 making it 16 then u would times 16 by its self 3 times like i showed u with the 2

OpenStudy (anonymous):

2*2=4..*2=8..*2=16..... 16*16=__..*16=(____) thats how u get ur answer

OpenStudy (anonymous):

oh thank you

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

i hope that helped

OpenStudy (anonymous):

yeah it did thanks

OpenStudy (anonymous):

any time

OpenStudy (anonymous):

i add you as a fan

OpenStudy (anonymous):

i added u as one too

OpenStudy (anonymous):

thanks kittykat

OpenStudy (anonymous):

ur welcome den